# Towards automated formal security analysis of SAML V2.0 Web Browser SSO standard — the POST/Artifact use case

### v1

@@ IV Security properties and analysis results
– property verified, – property falsified
TABLE II: Analysis results for each property and protocol variant.
The security properties can be roughly grouped into authentication properties, secrecy properties and freshness properties. We consider all variants of authentication properties described in the seminal paper [36] — namely weak, non-injective and injective agreement. Secrecy properties include secrecy of assertions and resources. Finally, we consider freshness properties of particular protocol parameters. Properties can be, furthermore, grouped by the participant from whose point of view we consider the protocol execution (client, SP and IdP), and by the exact assumptions on entities controlled by the adversary.
It is important to note that the desired security properties are not explicit or precise in the official SAML documentation. The said documentation includes the “Security considerations” non-normative document [37] containing high-level descriptions of the desired privacy and security properties as well as the threat model. Additionally, it gives multiple practical guidelines for implementing systems based on the SAML framework in the secure manner. In comparison with SAML documentation, we consider a more comprehensive set of security properties, and provide their precise formalization.
In total, we have considered 28 security properties and have either formally verified or falsified every property for each of the 8 protocol variants. The complete overview of the verification results is given in Table II, and the rest of this section gives an overview of the most relevant properties.
IV-A Client PoV
First, we investigate the security properties from the client’s point of view. That is, we assume that a client that has successfully completed a protocol session apparently with the SP $S$ and the IdP $I$, and investigate what guarantees hold for the client under different assumptions on honesty of other protocol principals. The honesty assumptions we consider are as follows:
@@ IV-A Client PoV
If a client $C$ completes a protocol session with the SP $S$ and the IdP $I$ with session parameters client ID $cid$, obtaining resource $res$ at URL $url$ then $C$ has registered that exact resource $res$ located at $url$ at the SP $S$.
@@
If a client $C$ registers a resource $res$ at the SP $S$ then that resource remains forever unknown to the adversary.
In the strong variant of property, we expect the secrecy to hold even if there exists some compromised IdP $I^{\prime}$ trusted by $SP$. This variant is included for sanity checking purposes as it should clearly fail — a compromised IdP $I^{\prime}$ trusted by $S$ can initiate a session, craft authentication responses for any client of its choice, and obtain their resources.
The formal analysis shows that the weaker version of the property holds for all protocols considered, whereas the strong version fails for all protocols variants as expected.
@@
If a client $C$ completes a protocol session with the SP $S$ and IdP $I$ with parameters client ID $cid$, and resource $res$ located at URL $url$ accompanied with temporary secret $ts$, then the temporary secret $ts$ remains forever unknown to the adversary.
Temporary secret $ts$ is of type fresh and is created and delivered in last step when SP delivers resource to client. This property means that if client manages to complete the protocol and to acquire resource in that concrete protocol run, it is not possible that the adversary obtained the resource compromising that particular session.
@@
If a client $C$ with ID $cid$ completes the protocol session with the SP $S$ obtaining resource $res$ accompanied with temporary secret $ts$, then the temporary secret $ts$ was created fresh after $C$ has requested the resource. This property holds unless $S$ is compromised.
@@
If a client $C$ with ID $cid$ completes the protocol session with the SP $S$ and IdP $I$ obtaining resource $res$ at $url$ with temporary secret $ts$, then $S$ has completed a protocol with $cid$ and the session parameters (the resource $res$, URL $url$ and temporary secret $ts$) all match.
This is a standard two party authentication property — if a client has $C$ completed a session with $S$ then there is a session of $S$ where the session parameters agree with those in the client session.
@@
If a client $C$ with id $cid$ completes the protocol session with the SP $S$ and IdP $I$ obtaining resource $res$ at URL $url$ with temporary secret $ts$ then $S$ has completed a protocol with client represented by $cid$ and the session parameters match and, additionally, there must not exist a client session of with same parameters $res$ and $cid$.
@@ IV-B Service Provider PoV
It is important to note that these properties between the client and the SP are not redundant or implied by the analogous properties from the PoV of the client. For example, SP – client resource secrecy property says that the temporary secret $ts$ served by the SP $S$ in some session remains unknown to the adversary as long as the client that registered the corresponding resource and the IdP $I^{\prime}$ used in that particular session are not compromised. This property guarantees that the honest client’s resources remain secret even in the scenarios when the client mistakenly establishes or completes sessions with compromised service providers — and these scenarios are not covered by the analogous properties from the PoV of the client. We do not include the SP – client injective agreement in the list of properties since it is trivially implied by the non-injective agreement — the session parameters (e.g., the ID $ida$ and the temporary secret $ts$ are generated fresh by $S$ making the SP’s session trivially unique).
@@
Next, we consider authentication properties between the SP and the IdP (namely the weak agreement and the non-injective agreement), and show that the weak and strong versions of both properties hold for all protocol variants. The agreement between the SP and the IdP is probably the most important property of the protocol as it captures the purpose of the SSO — if a SP accepts the assertion from an IdP, then IdP must have issued the assertion after authenticating the client. In all properties between SP and IdP we make no assumptions on the honesty of the clients. Hence, the properties hold when all clients are under the control of the adversary.
In the following property descriptions, $ida$ is the unique identifier of an <AuthnRequest> message while $mh$ is an unique message handle of an <Artifact> and $idr$ is unique identifier of an <ArtifactResolve> message.
@@
If a SP $S$ completes a protocol with client with ID $cid$ and the IdP $I$, then $I$ had completed a protocol and authenticated the client with ID $cid$. Furthermore, and all other parameters ($mh$, $ida$ and $idr$) match.
@@
In addition to the non-injective agreement above, this property requires that the authentication of the client with ID $cid$ is fresh in sense that it happened after the $S$ has sent the authentication request.
The analysis verifies this properties for 6 protocol variants and falsifies the property for 2 protocol variants. The property does not hold in variants where the authentication requests are not signed and the IDs are cryptographically weak. In the attack that violates this property the adversary guesses the ID used for authentication request and uses it to construct an authentication request that appears to be originating from $S$. The result is that the client authentication is triggered before the genuine authentication request is issued by the SP $S$. We do not consider the violation of this property to be a serious drawback of the protocol in practice as discussed in Section V.
@@
If a SP $S$ accepts an assertion containing an assertion secret $sec$, then that assertion remains unknown to the adversary.
@@ IV-C Identity provider POV
If an IdP $I$ has completed the protocol and issued an assertion to the SP $S$ in response to an <AuthnRequest> with ID $ida$, then $S$ has issued the <AuthnRequest> with ID $ida$ and has chosen $I$ as the IdP for the session.
Formal analysis verifies this properties for the four protocol variants where the authentication request are required to be signed and falsifies the property for the four protocol variants where the authentication requests are unsigned.
@@
In the attack falsifying the property, adversary acting as a client obtains a genuine <AuthnRequest> from the SP $S$, but replaces the identifier $ida$ with a different value before forwarding the request to the $I$. The $I$ will, now, complete the protocol — it will resolve the artifact even though there is a mismatch between the $ida$ parameters in the sessions of $I$ and $S$.
Note that the SP $S$ will not complete the protocol session — in the last step it will reject the assertion since the “InResponseTo” field in the assertion does not match the original <AuthnRequest> ID $ida$. This behavior agrees with the assertion processing rules in the SAML specifications. Again, we do not consider this failure to have a significant impact in real world deployments.
@@
If an IdP $I$ has completed the protocol and issued an assertion to $S$, then $S$ has issued the corresponding <AuthnRequest> choosing $I$ as the identity provider and also requested the <Artifact> that $I$ resolved. Hence $I$ and $S$ agree on the parameters of the <AuthnRequest> and the <Artifact>.
Formal analysis falsifies this property for every protocol variant. This is somewhat surprising at the first glance, but deeper analysis reveals that the property should indeed fail. The fundamental reason is similar to non-injective pre-agreement above — from the IdPs point of view there is no way of verifying that the artifact the SP is attempting to resolve corresponds to the original authorization request.
In the protocol execution falsifying this property, adversary takes the role of a compromised client and perofrms what we call an artifact-session confusion attack as follows:
@@
The end result of the execution is that IdP has completed a protocol session but there is no session of $S$ that agrees on the parameters of <AuthnRequest> and <Artifact>.
This attack is consistent with SAML specification for the following reasons: First, there is no way for $S$ to check that the <Artifact> received in step 4 corresponds to a different session – the artifact is opaque by design and the only contains the information about its originator $I$. Second, it is neither required or possible for $I$ to perform additional checks during the artifact resolution protocol in step 5. We were also able to confirm this attack is possible in practice by launching it against our simple deployment consiting of Shibboleth [33] Service Provider v3 and Identity Provider v4.
Again, the damage to the protocol is limited due to the fact that the a SAML compliant SP $S$ will not complete the protocol and issue the resource since abort the protocol after detecting the mismatch between the received assertion and the original <AuthnRequest>. Hence, we believe that this attack has no direct impact on the security of deployments, but it highlights the necessity of following the assertion processing rules specified in the SAML standard.
@@
IV-D Challenges of automated verification
The Tamarin meta model (with comments removed) contains 812 lines of code, about half of which correspond to protocol rules and half to lemmas and security properties. Each of the 8 concrete models for protocol variants is of roughly the same size as the meta model. Even though the models are of moderate size, the analysis proved challenging with many proofs never terminating, even in the simplified setting with a bounded number of sessions. Hence, several insights were needed to achieve termination and to finally verify/falsify all properties in a general setting with an unbounded number of sessions.
One source of complexity were the TLS channel rules initially implemented through facts — similar how TLS channels have been modelled in Tamarin before [35, 38] (which is, in turn, similar to confidential and authentic channels in the Tamarin documentation). This approach, however, resulted in exploding complexity during the precomputation phase due to the enumeration of all possible interactions of honest protocol participants and the adversary via the TLS channels. We addressed this complexity by, first, modelling TLS channels using a combination of private function symbols and adversary rules as described in Section III-B. This, along with a source lemma to resolve partial deconstructions, significantly reduced the complexity of precomputation (while, of course, moving some of the complexity to the main analysis phase). Secondly, we proved a number of TLS-related auxiliary lemmas further reducing the complexity of the analysis.
For most properties, further optimizations were needed. One other source of complexity is the large number of fresh values used in the protocol (in particular, most SAML messages require IDs which are modelled as fresh values). In each of our protocol models, there were roughly 50 refined sources for the $KU(\sim{}t)$ fact which translates to 50 new branches in the search tree each time a goal of such type is processed. We finally tamed the complexity by writing proof-guiding oracles for each problematic lemma, or groups of related lemmas.
Also, we significantly improved the analysis time of the security properties considered from the SP and the IdP PoV by removing all client rules from the model. Since there is no assumptions on the honesty of clients in those properties and, furthermore, all client actions can be performed by the adversary (as evidenced by the executable lemmas), there is no loss of generality here.
@@ 4 SP and IdP PoV properties 425s 431s 403s 403s 1569s 242s 253s 230s
TABLE III: Total analysis times (in seconds) for all lemma groups and protocol variants.
Final analysis was performed using Tamarin version 1.6.1 on an 8-core/16-thread desktop machine with Intel i9-9900K CPU and 32G of RAM. The total analysis time for all properties and protocol variants is close to 10 hours (160 hours of CPU time); the breakdown per protocol variant and the type of lemma is given in Table III.
V Summary and discussion of main results
Based on the results of the analysis we make the following claims on the POST/Artifact use case:
V-A The POST/Artifact use case satisfies desired security goals for a Web SSO system.
@@ V Summary and discussion of main results
The artifact-session confusion attack breaking the IdP-SP non-injective agreement property is interesting, but the extent of the attack is that an IdP can be made to resolve authentication responses that SP will never consume. Hence, we believe that there is no practical impact of this attack, but repeat that care must be taken by the SP to follow the assertion processing rules completely.
@@
V-B Details left unspecified by the SAML standard can have an impact on the exact security properties provided by the use case.
The protocols with cryptographically strong IDs have slightly different security properties then those with weak IDs. Admittedly, the difference is subtle, only present when combined with unsigned authentication requests, and unlikely to have any practical consequences. However, we believe the results validate the general concern that the seemingly purely functional choices left to developers by the SAML standard can end up having an impact to the security of the implemented system. However, the formal analysis did not discover any difference between the two different relay state mechanisms considered.
V-C There is no clear mapping between SAML optional protection mechanisms and provided security guarantees.
Adding signatures to authentication requests does not have a significant influence on the security of the use case. For the most important security properties there is no difference between the variants where signatures are required and when they are not. Besides the above mentioned authentication freshness issue, the only difference is the IdP-SP non-injective pre-agreement property where the adversary can make the IdP issue unsolicited authentication responses (that the SP will ultimatelly reject). This claim should be taken only in the limited context of symbolic protocol analysis — in a real-world deployment an IdP that requires signed authentication requests may be much less vulnerable to various low-level attacks (compared to an IdP that does not) simply due to a reduced attack surface as described in [37].
@@
Hence, it is hard for SAML users to decide what optional security mechanisms to include since it is highly non-trivial to deduce the impact of those security mechanisms on the final SSO system.
VI Related work
The most notable example of a formal analysis of a SAML SSO use case is the research of Armando et al. [18, 19, 20].
They model and analyze the Redirect/POST use case, as well as an actual deployment by Google, using a SAT-based model checker SATMC [39]. They consider two security properties — authentication based on non-injective agreement and the resource secrecy. The most significant contribution is the discovery of a crucial flaw in the Google’s deployment — the SP identifier was left out from the <AuthnRequest> message leading to an attack breaking both security properties considered. The authors also note that the tool fails to find similar attacks when the SP identifier is included in the message as required by the SAML specifications.
In the only other relevant attempt at a formal analysis of SAML use cases [40], Ferdous and Poet analyze the same use case as in [20], and create a simpler model, but one that includes an explicit user authentication step and a rudimentary digital identity model. The security properties are again similar to [20] and include basic authentication and secrecy properties. They use the High-Level Protocol Specification Language (HLPSL) to specify the protocols and the Automated Validation of Internet Security Protocols and Applications (AVISPA) tool with the On-the-fly Model-Checker (OFMC) backend for analysis. Besides the SAML use case, they also investigate other SSO protocols namely OpenID Connect and OAuth. The analysis by the tool discovered no attacks, but the authors do not make explicit claims on the exact setting in which the properties hold.

### v2

@@ V-B Overview of security properties and analysis results
   – property verified, – property falsified
Table III: Analysis results for each property and protocol variant.
V-C Client PoV
First, we investigate the security properties from the client’s point of view. That is, we assume that a client that has successfully completed a protocol session apparently with the SP $S$ and the IdP II, and investigate what guarantees hold for the client under different assumptions on honesty of other protocol principals. The honesty assumptions we consider are as follows:
@@ V-C Client PoV
If a client $C$ completes a protocol session with the SP $S$ and the IdP $I$ with session parameters client ID cidcid, obtaining resource resres at URL urlurl then $C$ has registered that exact resource resres located at urlurl at the SP $S$.
@@
If a client $C$ registers a resource resres at the SP $S$ then that resource remains forever unknown to the adversary.
As described in V-A the strong variant is included for sanity checking purposes, and it fails for all protocol variants, as expected.
The formal analysis shows that the weaker version of the property holds for all protocols considered
@@
If a client $C$ completes a protocol session with the SP $S$ and IdP $I$ with parameters client ID cidcid, and resource resres located at URL urlurl accompanied by temporary secret $ts$, then the temporary secret $ts$ remains forever unknown to the adversary.
Temporary secret $ts$ is of type fresh and is created and delivered in the last step when SP delivers resource to client. This property means that if the client manages to complete the protocol and to acquire a resource in that concrete protocol run, it is not possible that the adversary obtained the resource compromising that particular session.
@@
If a client $C$ with ID cidcid completes the protocol session with the SP $S$ obtaining resource resres accompanied by a temporary secret $ts$, then the temporary secret $ts$ was created fresh after $C$ has requested the resource. This property holds unless $S$ is compromised.
@@
If a client $C$ with ID cidcid completes the protocol session with the SP $S$ and IdP $I$ obtaining resource resres at urlurl with temporary secret $ts$, then $S$ has completed a protocol with cidcid and the session parameters (the resource resres, URL urlurl and temporary secret $ts$) all match.
This is a standard two party authentication property — if a client $C$ has completed a session with $S$ then there is a session of $S$ where the session parameters agree with those in the client session.
@@
If a client $C$ with ID cidcid completes the protocol session with the SP $S$ and IdP $I$ obtaining resource resres at URL urlurl with temporary secret $ts$ then $S$ has completed a protocol with client represented by cidcid and the session parameters match and, additionally, there must not exist a client session of with same parameters resres and cidcid.
@@ V-D Service Provider PoV
It is important to note that these properties between the client and the SP are not redundant or implied by the analogous properties from the PoV of the client. For example, SP – client resource secrecy property says that the temporary secret $ts$ served by the SP $S$ in some session remains unknown to the adversary as long as the client that registered the corresponding resource and the IdP $I^{\prime}$ used in that particular session are not compromised. This property guarantees that the honest client’s resources remain secret even in the scenarios when the client mistakenly establishes or completes sessions with compromised service providers — and these scenarios are not covered by the analogous properties from the PoV of the client. We do not include the SP – client injective agreement in the list of properties since it is trivially implied by the non-injective agreement — the session parameters (e.g., the ID idaida and the temporary secret $ts$ are generated fresh by $S$ making the SP’s session trivially unique).
@@
Next, we consider authentication properties between the SP and the IdP (namely the weak agreement and the non-injective agreement), and show that the weak and strong versions of both properties hold for all protocol variants. The agreement between the SP and the IdP is probably the most important property of the protocol as it captures the purpose of the SSO — if an SP accepts the assertion from an IdP, then IdP must have issued the assertion after authenticating the client. In all properties between SP and IdP we make no assumptions on the honesty of the clients. Hence, the properties hold when all clients are under the control of the adversary.
In the following property descriptions, idaida is the unique identifier of an <AuthnRequest> message while $mh$ is an unique message handle of an <Artifact> and idridr is unique identifier of an <ArtifactResolve> message.
@@
If an SP $S$ completes a protocol with the client with ID cidcid and the IdP $I$, then $I$ had completed a protocol and authenticated the client with ID cidcid. Furthermore, and all other parameters ($mh$, idaida and idridr) match.
@@
In addition to the non-injective agreement above, this property requires that the authentication of the client with ID cidcid is fresh, in the sense that it happened after the $S$ has sent the authentication request.
The analysis verifies these properties for 6 protocol variants and falsifies the property for 2 protocol variants. The property does not hold in variants where the authentication requests are not signed, and the IDs are cryptographically weak. In the attack that violates this property, the adversary guesses the ID used for the authentication request and uses it to construct an authentication request that appears to be originating from $S$. The result is that the client authentication is triggered before the genuine authentication request is issued by the SP $S$. We do not consider the violation of this property to be a serious drawback of the protocol in practice, as discussed in Section VI.
@@
If an SP $S$ accepts an assertion containing an assertion secret secsec, then that assertion remains unknown to the adversary.
@@ V-E Identity provider POV
If an IdP $I$ has completed the protocol and issued an assertion to the SP $S$ in response to an <AuthnRequest> with ID idaida, then $S$ has issued the <AuthnRequest> with ID idaida and has chosen $I$ as the IdP for the session.
Formal analysis verifies these properties for the four protocol variants where the authentication request are required to be signed and falsifies the property for the four protocol variants where the authentication requests are unsigned.
@@
In the attack falsifying the property, the adversary acting as a client obtains a genuine <AuthnRequest> from the SP $S$, but replaces the identifier idaida with a different value before forwarding the request to the $I$. The $I$ will, now, complete the protocol — it will resolve the artifact even though there is a mismatch between the idaida parameters in the sessions of $I$ and $S$.
Note that the SP $S$ will not complete the protocol session — in the last step it will reject the assertion since the <InResponseTo> field in the assertion does not match the original <AuthnRequest> ID idaida. This behavior agrees with the assertion processing rules in the SAML specifications. Again, we do not consider this failure to have a significant impact in real world deployments.
@@
If an IdP $I$ has completed the protocol and issued an assertion to $S$, then $S$ has issued the corresponding <AuthnRequest> choosing $I$ as the identity provider and also requested the <Artifact> that $I$ resolved. Hence, $I$ and $S$ agree on the parameters of the <AuthnRequest> and the <Artifact>.
Formal analysis falsifies this property for every protocol variant. This is somewhat surprising at the first glance, but deeper analysis reveals that the property should indeed fail.
In the protocol execution falsifying this property, the adversary takes the role of a compromised client and performs what we call an artifact-session confusion attack as follows:
@@
The end result of the execution is that IdP has completed a protocol session, but there is no session of $S$ that agrees on the parameters of <AuthnRequest> and <Artifact>. This vulnerability allows an adversary participating in a second session to manipulate the Service Provider (SP) into resolving an artifact from the first session. It is important to clarify that this vulnerability does not elevate the adversary’s privileges to completely bypass authentication — the adversary must still authenticate with the Identity Provider (IdP) to obtain the artifact in the second session.
This attack is consistent with SAML specification for the following reasons: First, there is no way for $S$ to check that the <Artifact> received in step 4 corresponds to a different session – the artifact is opaque by design and the only contains the information about its originator $I$. Second, it is neither required or possible for $I$ to perform additional checks during the artifact resolution protocol in step 5. We were also able to confirm this attack is possible in practice by launching it against our simple deployment consisting of Shibboleth [44] Service Provider v3 and Identity Provider v4.
Again, the damage to the protocol is limited due to the fact that a SAML compliant SP $S$ will not complete the protocol and issue the resource — it will abort the protocol after detecting the mismatch between the received assertion and the original <AuthnRequest>. Specifically, the standard states that the SP must verify that the <InResponseTo> attribute in the bearer <SubjectConfirmationData> equals the ID of its original <AuthnRequest> message, unless the response is unsolicited [22] — and our model implements this. Hence, we believe that this attack has no direct impact on the security of deployments, but it highlights the necessity of following the assertion processing rules specified in the SAML standard.
@@
V-F Challenges of automated verification
@@ 4 SP and IdP PoV properties 425s 431s 403s 403s 1569s 242s 253s 230s
Table IV: Total analysis times (in seconds) for all lemma groups and protocol variants.
The Tamarin meta model (with comments removed) contains 812 lines of code, about half of which correspond to protocol rules and half to lemmas and security properties. Each of the 8 concrete models for protocol variants is of roughly the same size as the meta model. Even though the models are of moderate size, the analysis proved challenging with many proofs never terminating, even in the simplified setting with a bounded number of sessions. Hence, several insights were needed to achieve termination and to finally verify/falsify all properties in a general setting with an unbounded number of agents and sessions.
Final analysis was performed using Tamarin version 1.6.1 on an 8-core/16-thread desktop machine with Intel i9-9900K CPU and 32G of RAM. The total analysis time for all properties and protocol variants is close to 10 hours (160 hours of CPU time); the breakdown per protocol variant and the type of lemma is given in Table IV.
The subsequent subsections elaborate on the key lessons learned that are essential for achieving finite-state verification.
V-F1 Combinatorial Explosion
The Tamarin prover exhaustively explores all possible rule applications within a protocol to reach the desired pre-conditions.
One source of complexity were the TLS channel rules initially implemented through facts — similar how TLS channels have been modelled in Tamarin before [36, 37] (which is, in turn, similar to confidential and authentic channels in the Tamarin documentation). This approach, however, resulted in exploding complexity during the precomputation phase due to the enumeration of all possible interactions of honest protocol participants and the adversary via the TLS channels. We addressed this complexity by, first, modelling TLS channels using a combination of private function symbols and adversary rules as described in Section IV-B. This, along with a source lemma to resolve partial deconstructions, significantly reduced the complexity of precomputation (while, of course, moving some of the complexity to the main analysis phase). Secondly, we proved a number of TLS-related helper lemmas, further reducing the complexity of the analysis.
Another source of complexity is the large number of fresh values used in the protocol —most SAML messages require IDs which are modelled as fresh values. In each of our protocol models, there were roughly 50 refined sources for the $KU(\sim{}t)$ fact which translates to 50 new branches in the search tree each time a goal of such type is processed. We finally tamed the complexity by writing proof-guiding oracles for each problematic lemma, or groups of related lemmas.
V-F2 Protocol and property abstraction
To achieve finite-state verification, security properties and/or protocol rules can be strategically and carefully abstracted. This reduces the complexity of the verification process, enabling it to complete within a reasonable time-frame.
Also, we significantly improved the analysis time of the security properties considered from the SP’s and the IdP’s PoV by removing all client rules from the model. Since there is no assumptions on the honesty of clients in those properties and, furthermore, all client actions can be performed by the adversary (as evidenced by the executable lemmas), there is no loss of generality here.
VI Summary and discussion of main results
Based on the results of the analysis, we make the following claims on the POST/Artifact use case:
VI-A The POST/Artifact use case satisfies desired security goals for a Web SSO system.
@@ VI Summary and discussion of main results
VI-B Details left unspecified by the SAML standard can have an impact on the exact security properties provided by the use case.
The protocols with cryptographically strong IDs have slightly different security properties than those with weak IDs. Admittedly, the difference is subtle, only present when combined with unsigned authentication requests, and unlikely to have any practical consequences. However, we believe the results validate the general concern that the seemingly purely functional choices left to developers by the SAML standard can end up having an impact to the security of the implemented system. However, the formal analysis did not discover any difference between the two different relay state mechanisms considered.
VI-C There is no clear mapping between SAML optional protection mechanisms and provided security guarantees.
Adding signatures to authentication requests does not have a significant influence on the security of the use case. For the most important security properties, there is no difference between the variants where signatures are required and when they are not. Besides the above-mentioned authentication freshness issue, the only difference is the IdP-SP non-injective pre-agreement property, where the adversary can make the IdP issue unsolicited authentication responses (that the SP will ultimately reject). This claim should be taken only in the limited context of symbolic protocol analysis — in a real-world deployment an IdP that requires signed authentication requests may be much less vulnerable to various low-level attacks (compared to an IdP that does not) simply due to a reduced attack surface as described in [23].
@@
Hence, it is hard for SAML deployers to decide what optional security mechanisms to include, since it is highly non-trivial to deduce the impact of those security mechanisms on the final SSO system.
VI-D Artifact-session confusion attack mitigation
The artifact-session confusion attack breaking the IdP-SP non-injective agreement property is interesting, but the extent of the attack is that an IdP can be made to resolve authentication responses that standards compliant SP will never consume. Hence, we believe that there is no practical impact of this attack, but repeat that care must be taken by the SP to follow the assertion processing rules completely. However, we do propose a modification to the artifact creation process defined within the SAML standard [29] that will mitigate the attack.
The existing standard mandates that artifacts be resolvable to their originating sender and exhibit single-use semantics. While SAML standard says that “Different types can be defined and used without affecting the binding” (Section 3.6.4 in [38]), it currently defines a single artifact format consisting of a random message handle and the <SourceId> uniquely identifying the “artifact issuer identity and the set of possible resolution endpoints“. The artifact receiver uses the <SourceId> to decide where to resolve the artifact. Note that <SourceId> processing plays a significant role for the security of the use-case — we discovered a man-in-the-middle attacks in the early version of the formal model where the <SourceId> was missing from the artifact.
Our proposed mitigation for the artifact-session confusion attack is to extend the artifact format and include a reference to the original SAML request within the artifact. Specifically for this use case, the artifact would contain the ‘id_authnrequest‘ identifier originating from the SAML Authentication Request sent by the SP to the IdP.
This approach strengthens session integrity by enabling the SP and IdP to verify the consistency between the received artifact and the corresponding SAML request. Even if the IdP receives an artifact containing a mismatch between the message handle and the ‘id_authnrequest‘ identifier, the <ArtifactResolve> message can be discarded, effectively thwarting artifact-session confusion attacks.
We formally model this proposed fix and formally verify that it does indeed solve the problem. We skip the details in the paper, but in [31] we make available the protocol model with the proposed fix and its complete analysis. These results demonstrate that ”IdP – SP non-injective pre-agreement,” ”IdP – SP non-injective agreement,” and ”IdP – SP injective agreement” properties all hold true for in all eight variants with the proposed fix.